How can you find the slope-intercept form of the line perpendicular to y=-3x, passing through the point (o,-3)?
Answer by THANApHD(104)

iF U'r graph is perpendicular to the y=-3x then Your graph will be
Y=3x+c, where c is the y intercept.
Use the cordinates to dtermine the c
-3=3*0+c
-3=0+c
so C willl be -3
So U'r function is
Y=3x-3